OpenStudy (anonymous):

$5,000 is invested in two different accounts yielding 3% and 3.5% interest. The interest earned on the two accounts is $155. Let x and y represent the two amounts invested. Complete the equation. = 155

OpenStudy (anonymous):

get a math book stupid

OpenStudy (anonymous):

or you could always just re read the question and figure its much simpler that this

OpenStudy (anonymous):

X=? y=? can you tell me what those are?

OpenStudy (anonymous):

i suck at algerbra so much

OpenStudy (anonymous):

they are both 5,000

OpenStudy (anonymous):

Let x and y represent the two amounts invested

OpenStudy (anonymous):

wait i read that wrong

OpenStudy (anonymous):

disregard that .03X+.035Y=155 -.03X -.03X .035Y=155-.03X \[\frac{ .035Y=155-.03X }{ .035 }\] \[Y=\frac{ 155-.03X }{ .035 }\]

OpenStudy (anonymous):

Solve for X ^

OpenStudy (anonymous):

.03X+.035Y=155 -.035Y -.035 .03X=155-.035Y \[\frac{ .03X=155-.035Y }{ .03 }\] \[X=\frac{ 155-.035Y }{ .03 }\] Solve for Y^

OpenStudy (anonymous):

backwards solve for the opposite of what i told you sorry

OpenStudy (anonymous):

X+Y=155

OpenStudy (anonymous):

he said thanks

OpenStudy (anonymous):

No Problem